7x-my = 1
7y-mx = 1
which value of "m" will make the equation have one solution, and how many values of "m" can make it possible.
Also have to find which value for "m" will make the equation have infinitely solutions.

Answer by Alan3354(69443)

7x-my = 1
7y-mx = 1
which value of "m" will make the equation have one solution, and how many values of "m" can make it possible.
Also have to find which value for "m" will make the equation have infinitely solutions.
---------
7x-my = 1
7y-mx = 1
-----
Find the 2 slopes.
7x-my = 1
my = 7x - 1
y = (7/m)*x - 1/m
slope = 7/m
---
7y-mx = 1
7y = mx + 1
y = (m/7)*x + 1/7
slope = m/7
===============
The slopes are different for all values of m except m = 7.
--> 1 solution.
========================
Also have to find which value for "m" will make the equation have infinitely solutions.
====
Infinite # of solutions --> coincident lines:
7x-my = 7y-mx
7x - 7y = my - mx
7(x-y) = m(y-x)
m = -7